OpenStudy (anonymous):

The state agriculture department asked random samples of Indiana farmers in each county whether they favored a mandatory corn check off program to pay for corn product marketing and research. In Tippecanoe County, 263 farmers were in favor of the program and 252 were not. In neighboring Benton County, 260 were in favor and 377 were not.

OpenStudy (anonymous):

b. Compute a 95% confidence interval for the difference between the proportions of farmers favoring the program in Tippecanoe County and in Benton County. Do you think opinions differed in the two counties?

OpenStudy (anonymous):

@reemii ?!?!?!

OpenStudy (reemii):

Do you know a statistic that is helpful for this? What kind of testing are we doing?

OpenStudy (anonymous):

95% confidence intervl?

OpenStudy (reemii):

What statistic are you going to use to compute the p-value?

OpenStudy (anonymous):

idk?

OpenStudy (reemii):

Have you already done testing for "comparison between proportions" ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

is ths is formula? : (p1hat-p2hat) plus/minus z* sqrt (p1hat (1-p1hat)/n1 plus (p2hat (1-p2hat/n2)

OpenStudy (reemii):

idk, i'm reading on some document that a statistic to use is: \( \frac{f_1-f_2}{S} \sim \mathcal N(0,1) \) where \(f_1,f_2\) are the appriximations of the proportions: T:(263+, 252-) -> estimation of the proportion: 263/(263+252)=\(f_1\) B:(260+, 377-) -> estimation of the proportion: 260/(260+377)=\(f_2\) and

OpenStudy (reemii):

\(S=\sqrt{f(1-f)}\sqrt{\frac 1{n_1}+\frac 1{n_2}} \) where \(f=\frac{n_1f_1+n_2f_2}{n_1+n_2}\).

OpenStudy (anonymous):

ok can u help me solve for the 95 percent confidence int

OpenStudy (reemii):

I can do something about rejecting or not the hypothesis that the means are the same. \(\frac{f_1-f_2}{S}\in [-z_\alpha,z_\alpha]\) with probability 95% , with \(z_\alpha=...\) oops i don't have the table..

OpenStudy (reemii):

aw.. Is this your statistic? \((\hat p_1-\hat p_2) \pm z \sqrt{ ( \hat p_1(1-\hat p_1)/n_1 + (\hat p_2 (1-\hat p_2/n_2)}\)

OpenStudy (anonymous):

i think yea

OpenStudy (reemii):

oops \((\hat p_2(1-\hat p_2))/n_2 \) <- there's an error in the big formula.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do i do

OpenStudy (reemii):

it is possible that my stat is the same as your stat (im too tired to compute now). Fine, when you know which stat you must use you're almost at the end.

OpenStudy (reemii):

you must choose the value of "\(\alpha\)". If you want a 99% confidence interval, choose \(\alpha=0.005\) (half of 1-0.99) and take a look in the table for the value of the quantile \(z_{\alpha}\).

OpenStudy (reemii):

then compute everything + replace \(z_\alpha\) by what you read in the table of quantiles of the normal(0,1) distribution. is it ok?

OpenStudy (anonymous):

umm i dont really understand.

OpenStudy (anonymous):

can u work it out plzzz? lol

OpenStudy (reemii):

Look at my old msg: \((f_1-f_2)/S \sim \mathcal N\). This means, the quantity has a probability of 99% to be contained in the interval \([z_{0.005},z_{1-0.005}]\). Ok for that?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

ok then wat

OpenStudy (anonymous):

?

OpenStudy (reemii):

from \((f_1-f_2)/S \in [z_1,z_2]\) we get to the interval \((f_1-f_2) \in [Sz_1,Sz_2]\) and then to the interval \([(f_1-f_2)+Sz_1, (f_1-f_2)+Sz_2]\). In your formula, \(z=z_{0.995}\), and this interval becomes \([(f_1-f_2)\pm zS]\) because \(z_{0.005}=-z_{0.995}\).

OpenStudy (reemii):

it's not a perfect explanation, i hope it helps.

OpenStudy (anonymous):

but do u know what the final answer will be?

OpenStudy (reemii):

no, you have to compute. take your stat and use the statement: \(n_1=263+252\), \(n_2=260+377\), \(\hat p_1=263/n_1\) etc.. and \(z_{0.995}\) must be looked up in a table.

OpenStudy (anonymous):

n1= 515, n2=637

OpenStudy (anonymous):

soo

OpenStudy (anonymous):

?

OpenStudy (reemii):

have you done something?

OpenStudy (anonymous):

no im not sure what to do

OpenStudy (anonymous):

I just need the conf interval!

OpenStudy (reemii):

I made a little mistake. you want a 95% conf.int. so \(\alpha=0.975\). Then \(z_{\alpha}=1.96\). Look, you need to replace everything in your stat: \((\hat p_1−\hat p_2)±z\sqrt{(\hat p_1(1−\hat p_1)/n_1+(\hat p_2(1−\hat p_2/n_2)}\). \(n_1=515, n_2=637, \hat p_1 = 262/n_1 = 0.51.., \hat p_2=0.408..\). Therefore, the conf.int. is \[ [0.00637.., 0.199] \] Try it and tell me if you understand.

OpenStudy (reemii):

oops

OpenStudy (anonymous):

so the official conf interval is (0.00637,0.199]

OpenStudy (anonymous):

(0.006, 0.199)<---

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

wait huh!

OpenStudy (anonymous):

is it (0.006, 0.199) or ?

OpenStudy (reemii):

it is rather [0.045, 0.16]

OpenStudy (anonymous):

[0.045, 0.16] is the answer?

OpenStudy (reemii):

did yo understand how we got there?

OpenStudy (anonymous):

so (0.05,0.16) is the answer?!??!

OpenStudy (anonymous):

so (0.05,0.16) is the answer? ru positive

OpenStudy (reemii):

using your formula, I obtain that indeed.

OpenStudy (anonymous):

ok thanks for ur help. and idk if thats the fomrula i just thought..

OpenStudy (reemii):

yw, gl